# Sharpness-Aware Minimization Can Hallucinate Minimizers

### v1

@@ 4.1 SAM can converge to hallucinated minimizers
4.2 Switching strategy for avoiding hallucinated minimizers
One obvious approach to avoiding hallucinated minimizers is to use a small perturbation radius $\rho>0$. Indeed, hallucinated minimizers do not arise when $\rho=0$ (i.e., when SAM is not applied). However, the perturbation itself is the key mechanism that regularizes against sharpness, making it desirable to use a moderately large value of $\rho$.
In this subsection, we introduce a simple yet effective heuristic for avoiding hallucinated minimizers, which we call switching. The idea is to use plain gradient descent for the first 10% of training iterations and then switch to SAM.
As shown in Figure 5(b), the switching strategy consistently drives the training loss to zero across all tested perturbation radii, including large values of $\rho$ for which standard SAM fails to converge to the true minima. Figure 6 further demonstrates the improved test accuracy achieved under the switching strategy. Not only does switching yield higher test accuracy, but it also reduces sensitivity to the choice of $\rho$. This is a notable advantage, since the perturbation radius $\rho$ is a difficult hyperparameter to tune in SAM.
The effectiveness of switching can be understood within our theoretical framework. The proof of Theorem 2.1 suggests that hallucinated minimizers tend to arise “near” local maximizers, and hence are more likely to occur in high-loss regions. This further implies that during the early stages of training, when loss values are large, SAM is particularly at risk of entering a basin of attraction of hallucinated minimizers. By first applying gradient descent to escape from high-loss regions before switching to SAM, this risk is mitigated. Consequently, our simple remedy ensures that SAM remains stable even for large perturbation radii.
A related mechanism was studied by Zhou et al. (2025), who applied SAM only in the later stages of training to improve generalization. In contrast, our results show that the switching strategy specifically prevents convergence to hallucinated minimizers, providing a complementary explanation for its effectiveness.
@@ 5 Conclusion
In this work, we identify a previously unrecognized failure mode of SAM: its tendency to converge to hallucinated minimizers. Our theoretical analysis establishes the existence of such minimizers under practical assumptions, and our empirical results validate the theory by demonstrating that SAM can indeed converge to them. To address this, we present a simple switching strategy that effectively avoids hallucinated minimizers.
Although our theoretical and empirical findings are consistent, gaps remain between the theoretical characterizations and broader empirical findings. These gaps open several interesting avenues for follow-up work.
One direction is to extend our theoretical analysis to the setting where SAM employs stochastic gradients rather than full-batch gradients. While our experiments suggest that hallucinated minimizers also arise in the stochastic case, a more rigorous theoretical understanding is desirable. Another direction is to analyze, from a theoretical standpoint, how likely it is for SAM to converge to hallucinated minimizers. Our experiments show that convergence to such minimizers is common, whereas our current theory only guarantees convergence within a local neighborhood of these points. Yet another direction is to extend the analysis to other variants of SAM. Our current results rely on the normalization of the ascent direction, which renders the magnitude of $\nabla f(x)$ irrelevant in constructing hallucinated minimizers. For SAM variants that incorporate gradient magnitude in the ascent step, a modified analysis would be necessary.
Reproducibility statement
We have taken extensive measures to ensure reproducibility. Complete proofs of all theorems, together with detailed assumptions, are provided in Appendix A and Appendix B. Experimental setups, including datasets and hyperparameters, are described in Appendices C and D. The implementation of our main experiments is available through a repository at https://anonymous.4open.science/r/SAM-can-hallucinate-minimizers-4B82.
@@ Appendix A Omitted details for Section 2
We begin by showing that the hallucinated minimizers cannot exist when the loss function is convex. For ease of exposition, throughout the appendices we use the following notation: for $x\in\mathbb{R}^{d}$, let
$x^{+}:=x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}.$
@@
If the function $f:\mathbb{R}^{d}\to\mathbb{R}$ is convex, then there is no point $x\in\mathbb{R}^{d}$ such that $\nabla f\left(x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}\right)=0$.
@@
Let $x^{\bullet}$ be a local maximizer with $f(x^{\bullet})=M$, and let $C\subseteq U$ be a compact ball centered at $x^{\bullet}$ such that $f(x)<f(x^{\bullet})$ for all $x\in C$. The strict inequality follows from $x^{\bullet}$ being an isolated critical point. Define $m:=\max_{\partial C}f(x)<M$ and consider the preimage $f^{-1}([M-\varepsilon,M])$ where $0<\varepsilon<M-m$. Let $C_{\varepsilon}$ denote the connected component of this preimage containing $x^{\bullet}\in f^{-1}([M-\varepsilon,M])$.
@@
Then, $x_{h}$ must be on the boundary of $C_{\varepsilon}$, and thus $f(x_{h})=M-\varepsilon$. Furthermore, since $x^{\bullet}\neq x_{h}$ is the only critical point of $f$ in $C_{\varepsilon}$, the gradient at $x_{h}$ does not vanish. Consequently, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=M-\varepsilon\}$ is an embedded $C^{1}$ hypersurface near $x_{h}$. By shrinking $V$ if necessary, we may assume $V\cap\partial C_{\varepsilon}=\Sigma$. Thus, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma$. By the method of Lagrange multipliers, we obtain
$\nabla g(x_{h})=\lambda\nabla f(x_{h}).$
That is, there exists $\lambda>0$ such that
$2(x_{\star}-x_{h})=\lambda\nabla f(x_{h})$
by Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
@@
which implies that $x_{h}$ is a hallucinated minimizer. ∎
@@
Take $x\in\partial C_{\varepsilon}$. If $f(x)>M-\varepsilon$, continuity of $f$ implies that there exists $r>0$ such that $f(y)>M-\varepsilon$ for all $y$ in an open ball centered at $x$ with radius $r$. This contradicts the fact that $x$ is a boundary point of $C_{\varepsilon}$.
@@
Consider any feasible direction $d\in\mathbb{R}^{d}$ with $\langle\nabla f(x_{h}),d\rangle\geq 0$. Let $\gamma\colon(-\varepsilon,\varepsilon)\rightarrow\mathbb{R}^{d}$ be a smooth curve such that $\gamma(0)=x_{h}$, $\gamma^{\prime}(0)=d$, and $\gamma(t)\in V$. Then, $f(\gamma(t))\geq M-\varepsilon$ for sufficiently small $t>0$. Moreover,
@@
Taking the feasible direction $d=\nabla f(x_{h})$ yields
$2\langle x_{h}-x_{\star},\nabla f(x_{h})\rangle=-2\lambda\|\nabla f(x_{h})\|^{2}\leq 0.$
@@
Suppose $f$ has a locally Lipschitz gradient. Then, Theorem 2.1 remains valid even when $x_{\star}$ is a local minimizer of $f$.
@@
$r^{\star}:=\min\Big\{\frac{\delta}{1+\frac{2\rho L}{\|\nabla f(x_{h})\|}},r\Big\}.$
@@
A.3 Most modern neural networks are real-analytic: Further discussion
We now formalize the claim that the neural network $\theta\mapsto h_{\theta}(X)$ is real-analytic in its parameters $\theta$. By standard arguments, the following two lemmas establish that the final loss function
$f(\theta)=\sum_{i=1}^{N}\ell(h_{\theta}(X_{i}),Y_{i})$
is real-analytic under the assumptions on $h_{\theta}$ stated in Section 2.
Lemma A.7.
If $h_{\theta}\colon\mathbb{R}^{d}\to\mathbb{R}^{h}$ is a neural network with real-analytic activation functions, then $h_{\theta}$ is real-analytic as a function of $\theta$.
Proof.
Fix input data $x$ and let $z$ denote the hidden states, which depend on $x$ and $\theta$. Then, $z$, hence the entire network $h_{\theta}$, is a finite composition of the following real-analytic mappings:
  • (Layer normalization) $z\mapsto\gamma\odot\frac{z-\mu_{L}(z)}{\sqrt{\sigma_{L}(z)+\varepsilon}}+\beta$, where $\mu_{L}{(z)}$ and $\sigma_{L}(z)$ are the per-sample mean and variance;
  • (Batch normalization) $z\mapsto\gamma\odot\frac{z-\mu_{B}{(z)}}{\sqrt{\sigma_{B}(z)+\varepsilon}}+\beta$, where $\mu_{B}{(z)}$ and $\sigma_{B}(z)$ are the per-channel mean and variance;
  • (Activation) $z\mapsto\sigma(z)$, where $\sigma$ is a real-analytic activation function;
  • (Softmax) $z\mapsto\mu(z)$, where $\mu$ is the softmax function;
  • (Average pooling) $z\mapsto Az$, where $A$ is a linear averaging operator;
  • (Residual connection) $z\mapsto z+F_{\theta}(z)$, where $F$ is real-analytic in $\theta$;
  • (Convolution layer) $(z,W_{i},b_{i})\mapsto W_{i}*z+b_{i}$, where * denotes the discrete convolution operator;
  • (Dropout) $z\mapsto m\odot z$, where $m$ is a masking operator;
  • (Linear layer) $(z,A_{i},b_{i})\mapsto A_{i}z+b_{i}$; and
  • (Attention layer) $(z,W_{Q},W_{K},W_{V})\mapsto\mu\big(\frac{{Q}{K}^{\top}}{\sqrt{d_{\mathrm{attn}}}}\big){V}$, where $Q=zW_{Q}$, $K=zW_{K}$, $V=zW_{V}$, and $d_{\mathrm{attn}}$ is the size of the attention matrix $Q,K$.
Since the composition of real-analytic functions is real-analytic, it follows that $h_{\theta}$ is real-analytic. ∎
Finally, the next lemma establishes that $f$ is real-analytic. This follows directly from the fact that the composition of real-analytic functions is real-analytic; hence the proof is omitted.
Lemma A.8.
Let $h_{\theta}$ be a (finite) neural network constructed with linear layers, attention, convolution, layer normalization, and real-analytic activation functions (all commonly used activation functions except ReLU are real-analytic). Then, for a real-analytic loss function $\ell$,
$f(\theta):=\frac{1}{N}\sum^{N}_{i=1}\ell(h_{\theta}(X_{i}),Y_{i})$
is real-analytic as a function of $\theta$.
A.4 Full proof of Theorem 2.2
@@ A.3 Most modern neural networks are real-analytic: Further discussion
Suppose $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic and $X$ is a bounded local maximizer set of $f$ with some $\delta>0$. Then, there exists $\varepsilon>0$ with the following property: if $x$ is a critical point in the $\delta$-neighborhood of $X$ with $f(x)\geq f(X)-\varepsilon$, then $x\in X$.
@@
If $S=\emptyset$, then the theorem holds for any $\varepsilon>0$, and we are done. Assume instead that $S\neq\emptyset$. For each point $x\in X$, Lemma 2.3 guarantees the existence of an open neighborhood $U_{x}$, a constant $C_{x}>0$, and an exponent $q_{x}\in(0,1)$ such that
@@
This implies $f(y)=f(X)$, which is a contradiction. Hence, $S\subseteq C:=\mathcal{N}_{\delta}(X)\setminus\bigcup_{x\in X}U_{x}$. (Note that $C\neq\emptyset$ since $S\neq\emptyset$). Since $C$ is compact, define
@@
Since the only critical points in $C_{\varepsilon}$ are those in $X$, we have $\nabla f(x_{h})\neq 0$. Thus, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=f(X)-\varepsilon\}$ is an embedded smooth hypersurface near $x_{h}$. By possibly shrinking $V$, we may assume $V\cap\partial C_{\varepsilon}=\Sigma$. Therefore, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma$.
@@
$\nabla g(x_{h})=\lambda\nabla f(x_{h}),$
which implies that there exists $\lambda>0$ such that
@@
The positivity of $\lambda$ follows from the same reasoning as in Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
@@
which shows that $x_{h}$ is a hallucinated minimizer. ∎

### v2

@@ 4.1 SAM can converge to hallucinated minimizers
Larger networks: full-batch ResNet-18 on CIFAR.
To test whether the same mismatch arises beyond small smooth networks, we run full-batch SAM on ResNet-18 for CIFAR-10/100 across $\rho\in\{0.0,0.5,1.0,1.5,2.0\}$ (Appendix C.2). Although ReLU networks fall outside our smooth/analytic assumptions, the same shifted-vs-original mismatch signatures are visible. Table 1 reports final-step metrics and shows that for $\rho>0$, the SAM loss $f(x_{T}^{+})$ and the shifted gradient norm $\|\nabla f(x_{T}^{+})\|$ can be orders of magnitude smaller than their unshifted counterparts. The corresponding trajectories (Appendix C.2) plateau at strictly positive $f(x_{T})$ while $f(x_{T}^{+})$ remains small, consistent with the hallucinated-minimizer mechanism.
Stochastic training on CIFAR-100.
Finally, we examine the practical behavior under standard mini-batch training for CIFAR-100 with ResNet-18. Figure 5 shows that SAM-only training becomes increasingly unstable as $\rho$ grows, while a short SGD warm-start (next subsection) improves stability and reduces sensitivity to $\rho$. Since our theory is deterministic/full-batch, we make no formal claims in the stochastic regime; we report these results as an empirical observation that the large-$\rho$ degradation can also arise in practice. Additional details are provided in Appendix C.3.
4.2 A short SGD warm-start as a practical safeguard
A trivial way to reduce the risk of hallucinated minimizers is to keep $\rho$ very small, but this can weaken the sharpness-regularization effect that motivates SAM. Instead, we consider a minimal and widely compatible SGD warm-start schedule: we run the base optimizer (SGD) for a short initial phase, and then enable SAM for the remainder of training. We use a fixed warm-start fraction $10\%$ in all experiments and do not tune it. Warm-start schedules and related heuristics have been considered in prior work (e.g., Andriushchenko and Flammarion (2022)); our focus here is not to propose a new training recipe, but to show that a short warm-start is sufficient to mitigate the hallucinated-minimizer failure mode identified in Sections 23.
Empirical effect.
In the full-batch MNIST setting, Figure 4 shows that SAM-only training frequently plateaus at positive loss for larger radii, whereas the SGD warm-start reliably avoids these plateaus and drives training loss to (near-)zero across the same $\rho$ sweep. In the stochastic CIFAR-100 setting, Figure 5 similarly shows improved stability and reduced sensitivity to $\rho$.
Connection to our theory.
The existence proof in Theorem 2.2 constructs hallucinated minimizers via geometry near local maximizer regions, and Theorem 3.1 shows that such points can have basins of attraction under the practical SAM dynamics. Early in training, iterates typically lie in high-loss regions, where encountering these basins is more likely. A brief SGD warm-start moves the trajectory away from high-loss regions before the SAM perturbation is enabled, reducing the chance of entering a hallucinated-minimizer basin without requiring $\rho$ to be small.
@@ 5 Conclusion
We identified a previously under-emphasized mismatch in standard SAM: the practical update uses the shifted gradient $\nabla f(x^{+})$, rather than the surrogate gradient $\nabla f^{\mathrm{SAM}}(x)$. This mismatch enables a distinct large-$\rho$ failure mode in nonconvex landscapes: SAM can stall at hallucinated minimizers—local minimizers of $f^{\mathrm{SAM}}$ that are not stationary points of the original objective. We established geometric existence mechanisms (including extensions to local maximizer sets), showed that isolated hallucinated-minimizer components can be locally attracting under the discrete-time SAM iteration, and characterized how hallucinated minimizers can inherit manifold structure via local invertibility of the perturbation map. Empirically, our diagnostics on neural network objectives exhibit mismatch signatures consistent with this mechanism, and we report that a short initial SGD warm-start can serve as a simple safeguard that improves stability and reduces sensitivity to $\rho$.
An important direction for future work is to develop analogous guarantees for stochastic mini-batch training as well as for nonsmooth architectures. It would also be valuable to analyze whether and how SAM variants that modify the ascent step suppress or amplify hallucinated minimizers.
@@ Appendix A Omitted details for Section 2
We begin by showing that the hallucinated minimizers cannot exist when the function is convex.
@@
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be convex and differentiable, and fix $\rho>0$. Then for any $x\in\mathbb{R}^{d}$ with $\nabla f(x)\neq 0$, we have $\nabla f(x^{+})\neq 0$, where $x^{+}:=x+\rho\frac{\nabla f(x)}{\|\nabla f(x)\|}$.
@@
Let $x^{\bullet}$ be an isolated strict local maximizer with $f(x^{\bullet})=M$. Choose a closed ball $C\subseteq U$ centered at $x^{\bullet}$ such that $f(x)<f(x^{\bullet})$ for all $x\in C\setminus\{x^{\bullet}\}$, and $x^{\bullet}$ is the only critical point of $f$ in $C$. Define $m:=\max_{\partial C}f(x)<M$ and consider the preimage $f^{-1}([M-\varepsilon,M])$ where $0<\varepsilon<M-m$. Let $C_{\varepsilon}$ denote the connected component of this preimage containing $x^{\bullet}\in f^{-1}([M-\varepsilon,M])$.
@@
Then, $x_{h}$ must be on the boundary of $C_{\varepsilon}$, and thus $f(x_{h})=M-\varepsilon$. Furthermore, since $x^{\bullet}$ is the only critical point of $f$ in $C_{\varepsilon}$ and $x_{h}\neq x^{\bullet}$, we have $\nabla f(x_{h})\neq 0$. Consequently, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=M-\varepsilon\}$ is an embedded $C^{1}$ hypersurface near $x_{h}$. By shrinking $V$ if necessary, we may assume $V\cap\partial C_{\varepsilon}=\Sigma$. Thus, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma=\{x\in V:\ f(x)=M-\varepsilon\}$ near $x_{h}$. Since $\nabla f(x_{h})\neq 0$, $\Sigma$ is a regular level set and hence an embedded $C^{1}$ hypersurface. Because $x_{h}$ maximizes $g$ over $C_{\varepsilon}$ and lies on $\partial C_{\varepsilon}$, it is also a local maximizer of $g$ over $\Sigma$. By the method of Lagrange multipliers, there exists $\lambda\in\mathbb{R}$ such that
$\nabla g(x_{h})+\lambda\nabla f(x_{h})=0.$
Moreover, $\lambda>0$ by Lemma A.4. It follows that
$2(x_{\star}-x_{h})=\lambda\nabla f(x_{h}).$
Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
@@
which implies that $x_{h}$ is a hallucinated minimizer. This shows that for each $\varepsilon\in(0,M-m)$ there is a corresponding radius $\rho>0$ such that $x_{h}$ is a hallucinated minimizer. Since $\rho(\varepsilon)$ is strictly increasing and continuous in $\varepsilon$ by Lemma A.5, there exists an interval of radii $\rho$ as claimed. ∎
@@
Take $x\in\partial C_{\varepsilon}$. If $f(x)>M-\varepsilon$, continuity of $f$ implies that there exists $r>0$ such that $f(y)>M-\varepsilon$ for all $y$ in an open ball $B_{r}(x)$ centered at $x$ with radius $r$. Since $B_{r}(x)\subset f^{-1}([M-\varepsilon,M])$ is connected and intersects $C_{\varepsilon}$, maximality of the connected component implies $B_{r}(x)\subset C_{\varepsilon}$, so $x$ would be an interior point. This contradicts the fact that $x$ is a boundary point of $C_{\varepsilon}$.
@@
Take the direction $d:=\nabla f(x_{h})$, for which $\langle\nabla f(x_{h}),d\rangle=\|\nabla f(x_{h})\|^{2}>0$. Let $\gamma\colon(-\varepsilon,\varepsilon)\to\mathbb{R}^{d}$ be a smooth curve with $\gamma(0)=x_{h}$ and $\gamma^{\prime}(0)=d$. Then, $f(\gamma(t))\geq M-\varepsilon$ for sufficiently small $t>0$. Moreover,
@@
Since $d=\nabla f(x_{h})$,
$2\langle x_{h}-x_{\star},\nabla f(x_{h})\rangle=-\lambda\|\nabla f(x_{h})\|^{2}\leq 0.$
@@
Lemma A.5.
In the proof of Theorem 2.2, define
$\rho(\varepsilon):=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|.$
Then $\rho(\varepsilon)$ is continuous and strictly increasing for $0<\varepsilon<M-m$.
Proof.
Since $C_{\varepsilon}$ is compact (Lemma A.2), the maximum defining $\rho(\varepsilon)$ is attained. Moreover, $\|x-x_{\star}\|$ cannot attain its maximum in $\mathrm{int}(C_{\varepsilon})$. Therefore
$\rho(\varepsilon)=\max_{x\in\partial C_{\varepsilon}}\|x-x_{\star}\|.$
Strict monotonicity.
Fix $0<\varepsilon_{1}<\varepsilon_{2}<M-m$. By construction of $C_{\varepsilon}$ in the proof of Theorem 2.2, $C_{\varepsilon_{1}}\subset C_{\varepsilon_{2}}$. Let $x_{1}\in\partial C_{\varepsilon_{1}}$ attain $\rho(\varepsilon_{1})$. By Lemma A.3, $f\equiv M-\varepsilon$ on $\partial C_{\varepsilon}$, so $\partial C_{\varepsilon_{1}}\cap\partial C_{\varepsilon_{2}}=\varnothing$. Thus $x_{1}\in\mathrm{int}(C_{\varepsilon_{2}})$, and there exists $\delta>0$ such that $B(x_{1},\delta)\subset C_{\varepsilon_{2}}$. Let $v:=(x_{1}-x_{\star})/\|x_{1}-x_{\star}\|$ and choose any $t\in(0,\delta)$. Then $x_{1}+tv\in C_{\varepsilon_{2}}$ and
$\rho(\varepsilon_{2})\geq\|x_{1}+tv-x_{\star}\|\geq\|x_{1}-x_{\star}\|+t>\|x_{1}-x_{\star}\|=\rho(\varepsilon_{1}).$
Thus $\rho(\varepsilon)$ is strictly increasing on (0,Mm)(0,M-m).
Continuity.
Fix $\varepsilon_{0}\in(0,M-m)$. By construction, $x^{\bullet}\in\mathrm{int}(C_{\varepsilon_{0}})$, and $x^{\bullet}$ is the only critical point of $f$ in $C$. Hence $\nabla f(x)\neq 0$ for all $x\in\partial C_{\varepsilon_{0}}$. In particular, the level value $M-\varepsilon_{0}$ is regular on $\partial C_{\varepsilon_{0}}$, so $\partial C_{\varepsilon}$ forms a $C^{1}$ hypersurface that varies continuously with $\varepsilon$ in a neighborhood of $\varepsilon_{0}$ (by the implicit function theorem, applied locally and combined via a finite cover of the compact set $\partial C_{\varepsilon_{0}}$).
Since $C_{\varepsilon}$ is compact for all $\varepsilon\in(0,M-m)$ and
$\rho(\varepsilon)=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|=\max_{x\in\partial C_{\varepsilon}}\|x-x_{\star}\|,$
standard maximum-continuity arguments for continuous functions over continuously varying compact sets yield that $\rho(\varepsilon)$ is continuous at $\varepsilon_{0}$. As $\varepsilon_{0}$ was arbitrary, $\rho$ is continuous on (0,Mm)(0,M-m). ∎
@@
Suppose $f$ has a locally Lipschitz gradient. Then Theorem 2.2 remains valid even when $x_{\star}$ is a local minimizer of $f$.
@@
$r^{\star}:=\min\left\{\frac{\delta}{1+\frac{2\rho L}{\|\nabla f(x_{h})\|}},r\right\}.$
@@
A.3 Full proof of Theorem 2.4
Lemma A.8 (Łojasiewicz (1963)).
$|f(p)-f(x)|^{q}\leq c\|\nabla f(x)\|\quad\text{for all }x\in U.$
@@
Suppose $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is real-analytic and $X$ is a bounded local maximizer set of $f$ for some $\delta>0$. Then, there exists $\varepsilon>0$ with the following property: if $x$ is a critical point in the $\delta$-neighborhood of $X$ with $f(x)\geq f(X)-\varepsilon$, then $x\in X$.
@@
If $S=\emptyset$, then the theorem holds for any $\varepsilon>0$, and we are done. Assume instead that $S\neq\emptyset$. For each point $x\in X$, Lemma A.8 guarantees the existence of an open neighborhood $U_{x}$, a constant $C_{x}>0$, and an exponent $q_{x}\in(0,1)$ such that
@@
This implies $f(y)=f(X)$, which contradicts $X=\operatorname*{arg\,max}_{z\in\mathcal{N}_{\delta}(X)}f(z)$ because $y\notin X$. Hence, $S\subseteq C:=\mathcal{N}_{\delta}(X)\setminus\bigcup_{x\in X}U_{x}$. (Note that $C\neq\emptyset$ since $S\neq\emptyset$). Since $C$ is compact, define
@@
Then $x_{h}\in\partial C_{\varepsilon}$ (otherwise we could move slightly in the direction $x_{h}-x_{\star}$ while staying in $C_{\varepsilon}$ and increase $g$). Hence $f(x_{h})=f(X)-\varepsilon<f(X)$, so $x_{h}\notin X$, and Lemma A.9 implies $\nabla f(x_{h})\neq 0$. Thus, there exists an open neighborhood $V$ of $x_{h}$ such that $\Sigma:=\{x\in V:f(x)=f(X)-\varepsilon\}$ is an embedded smooth hypersurface near $x_{h}$. By possibly shrinking $V$, we may assume $V\cap\partial C_{\varepsilon}=\Sigma$. Therefore, maximizing $g(x)$ over $C_{\varepsilon}$ is locally equivalent to maximizing $g(x)$ over the hypersurface $\Sigma$.
@@
$\nabla g(x_{h})+\lambda\nabla f(x_{h})=0.$
Equivalently,
@@
Moreover, $\lambda>0$ follows from the same argument as in Lemma A.4. Taking norms of both sides yields $\lambda=\frac{2\rho}{\|\nabla f(x_{h})\|}$. Therefore,
@@
which shows that $x_{h}$ is a hallucinated minimizer. Moreover, just like in the proof of Theorem 2.2, the above argument applies for all
$0<\varepsilon<\min\{\varepsilon_{1},\varepsilon_{2}\}.$
Hence, by Lemma A.10 below, we conclude that the set of radii for which a hallucinated minimizer exists contains a nontrivial interval. ∎
Lemma A.10 (Monotonicity/continuity of $\rho(\varepsilon)$ for maximizer-set construction).
With $C_{\varepsilon}$ defined as the connected component of $f^{-1}([f(X)-\varepsilon,f(X)])$ containing $X$, the map $\rho(\varepsilon):=\max_{x\in C_{\varepsilon}}\|x-x_{\star}\|$ is continuous and strictly increasing for $\varepsilon\in(0,\min\{\varepsilon_{1},\varepsilon_{2}\})$.
Proof.
The proof is identical to Lemma A.5 and uses only compactness, nestedness $C_{\varepsilon_{1}}\subset C_{\varepsilon_{2}}$ for $\varepsilon_{1}<\varepsilon_{2}$, and connectedness. ∎
A.4 When neural networks are real-analytic: sufficient conditions
Theorem 2.4 assumes that the objective $f$ is real-analytic. This assumption is used only to invoke the Łojasiewicz inequality (Lemma A.8) and thereby rule out pathological accumulations of critical points near a local maximizer set. Here we record a simple sufficient condition showing that real-analyticity holds for a class of smooth (deterministic) neural network objectives. In particular, it covers the full-batch MNIST experiment with Tanh activations in Section 4.1.
Throughout, we consider a finite dataset $\{(X_{i},Y_{i})\}_{i=1}^{N}$ and a deterministic network map $\theta\mapsto h_{\theta}(X_{i})$ (i.e., we exclude stochastic layers such as dropout from this discussion). We also assume normalization layers use the standard numerical constant $\varepsilon_{\mathrm{norm}}>0$ to avoid division by zero.
Lemma A.11 (Real-analyticity of common smooth building blocks).
Fix $\varepsilon_{\mathrm{norm}}>0$. The following mappings are real-analytic on their natural domains:
  • affine/linear maps and convolutions;
  • residual additions and average pooling;
  • softmax and log-softmax;
  • layer normalization and batch normalization in the standard form
    $\mathrm{Norm}(z)=\gamma\odot\frac{z-\mu(z)}{\sqrt{\sigma(z)+\varepsilon_{\mathrm{norm}}}}+\beta,$
    where $\mu(z)$ and $\sigma(z)$ are (per-sample or per-batch) mean/variance functions of $z$.
Moreover, any elementwise real-analytic activation $\sigma(\cdot)$ (e.g., Tanh, sigmoid, softplus, GELU, swish) is real-analytic.
Proof.
Affine maps, convolutions, additions, and average pooling are polynomial/linear maps, hence real-analytic. Softmax and log-softmax are compositions of $\exp$, finite sums, divisions by strictly positive quantities, and $\log$ applied to strictly positive quantities, hence real-analytic.
For normalization, note that $\mu(z)$ and $\sigma(z)$ are polynomials in the entries of $z$. Since $\sigma(z)\geq 0$ and $\varepsilon_{\mathrm{norm}}>0$, the quantity $\sigma(z)+\varepsilon_{\mathrm{norm}}$ is strictly positive, and the map $t\mapsto 1/\sqrt{t}$ is real-analytic on $(0,\infty)$. Thus $\mathrm{Norm}(z)$ is a composition of real-analytic maps, hence real-analytic. Finally, elementwise real-analytic activations preserve real-analyticity under composition. ∎
Lemma A.12.
Assume that for each fixed input $X$, the network output $\theta\mapsto h_{\theta}(X)$ is obtained by composing the building blocks in Lemma A.11 with real-analytic activations. Let the loss $\ell(\cdot,Y)$ be real-analytic in its first argument for each label $Y$ (e.g., squared loss; cross-entropy combined with softmax). Then the finite-dataset objective
$f(\theta):=\frac{1}{N}\sum^{N}_{i=1}\ell(h_{\theta}(X_{i}),Y_{i})$
is real-analytic in $\theta$.
Proof.
By Lemma A.11 and closure of real-analytic functions under composition, each map $\theta\mapsto h_{\theta}(X_{i})$ is real-analytic. Composing with the real-analytic loss $\ell(\cdot,Y_{i})$ preserves real-analyticity, and finite sums preserve real-analyticity. Hence $f$ is real-analytic. ∎